OpenStudy (anonymous):

How do I find the arc measurement of a sector with the area 38.485 ft2 and radius 7?

OpenStudy (anonymous):

If the radius is 7, the entire circle would have area pi*7^2 = 49*pi =~ 153.938 But we only have 38.485, so divide that by (49*pi) and we have: 38.485 / (49*pi) =~ 0.25, that's almost exactly 1/4 of a circle (90 degrees or pi/2 radians) *Taken from a different source

OpenStudy (anonymous):

Why would you divide it by (49π)? @EducateMe

OpenStudy (anonymous):

Because the whole complete circle is 49pi. We don't have a complete circle though, so we divide by 49pi. In other words, we are chopping up the circle into sectors

OpenStudy (anonymous):

Ohh okay. Thank you! c: